### v1

# MIRAGE: KV Cache Optimization through Parameter Remapping for Multi-tenant LLM Serving

Abstract.
KV cache accelerates LLM inference by avoiding redundant computation, at the expense of memory. To support larger KV caches, prior work extends GPU memory with CPU memory via CPU-offloading. This involves swapping KV cache between GPU and CPU memory. However, because the cache updates dynamically, such swapping incurs high CPU memory traffic. We make a key observation that model parameters remain constant during runtime, unlike the dynamically updated KV cache. Building on this, we introduce MIRAGE, which avoids KV cache swapping by remapping, and thereby repurposing, the memory allocated to model parameters for KV cache. This parameter remapping is especially beneficial in multi-tenant environments, where the memory used for the parameters of the inactive models can be more aggressively reclaimed. Exploiting the high CPU–GPU bandwidth offered by the modern hardware, such as the NVIDIA Grace Hopper Superchip, we show that MIRAGE significantly outperforms state-of-the-art solutions, achieving a reduction of 44.8%-82.5% in tail time-between-token latency, 20.7%-99.3% in tail time-to-first-token latency, and 6.6%-86.7% higher throughput compared to vLLM.
copyright: none
1. Introduction
The memory demands of Large Language Models (LLMs) are growing at a much faster rate than GPU memory capacity, making memory a major bottleneck for efficient inference serving (Hanindhito et al., 2024; Kwon et al., 2023; Touvron et al., 2023b; Dubey et al., 2024; Nvidia, 2025a, d, c). KV cache (Pope et al., 2023; Kwon et al., 2023) was introduced as a significant way to reduce the inference latency: it stores intermediate results from previously generated tokens to avoid recomputing tensor values. To harness the impact of KV cache on the performance of LLM inference serving, numerous memory management techniques are proposed (Pope et al., 2023; Lee et al., 2024a; Liu et al., 2024a; Prabhu et al., 2025; Jin et al., 2024; Adnan et al., 2024), including paging-based mechanisms for runtime KV cache management (Kwon et al., 2023). However, despite these optimizations, when the KV cache size hits the limit of the GPU memory, one must recompute key-value pairs, increasing serving latency (Kwon et al., 2023) (Figure 1). The KV-cache bottleneck increases further in multi-tenant environments where multiple LLMs share and contend over the GPU(s) resources (Yu et al., 2025).
Using CPU memory as an extension of the GPU memory (also called CPU-offloading) is one way to ease the pressure on GPU memory (Peng et al., 2020; Huang et al., 2020; Rajbhandari et al., 2021; Ren et al., 2021; Lim et al., 2021; Sheng et al., 2023; Jiang et al., 2024; Xu et al., 2024). However, it comes at the cost of increased latency or reduced throughput, depending on the CPU-GPU transfer bandwidth. Prior work (Jiang et al., 2024) explored offloading portions of attention computation to CPUs, allowing KV cache to be generated and stored in CPU memory rather than occupying GPU memory. This allowed the CPU–GPU system to handle larger batch sizes; however, limited CPU performance in attention computation can become a bottleneck, potentially increasing token generation latency by about 10-20% in certain scenarios (Jiang et al., 2024).
To make it practical, new advancements from the hardware community have dramatically increased the CPU-GPU data transfer bandwidth: the latest Grace Hopper superchip (GH200) (Nvidia, 2025c) introduces a new NVLink-C2C interface between CPU and GPU, offering 7×7$\times$7 × bandwidth of the H100 (Nvidia, 2025d) and 14×14$\times$14 × that of the A100 (Nvidia, 2025a) (Figure 1). This increased CPU-GPU bandwidth opens new opportunities to design low-latency and high-throughput LLM inference serving systems. At the same time, with the availability of higher bandwidth, the limited parallelism of CPUs, rather than the bandwidth, may emerge as the new performance bottleneck (§ 3.1).
(a) Existing LLM systems.
(b) GPU Trends.
Figure 1. (a) Existing LLM systems perform recomputation when KV cache is exhausted, leading to a significant increase in latency under high request rates (OPT-13b with ShareGPT). (b) CPU-GPU bandwidth in Nvidia GPUs has increased significantly since GH200, bringing CPU-offloading opportunities.
Existing work, such as Pie (Xu et al., 2024), leverages this extra CPU-GPU bandwidth by offloading the KV cache of future layers to CPU memory and swapping back to GPU memory as required, thereby increasing the effective KV cache size (see Figure 2 (left side)). This allows for larger batches or longer sequences without triggering recomputation, thereby reducing inference serving latency and improving serving throughput. However, the frequent updates to the KV cache in GPU memory require synchronization during KV cache swapping, introducing overheads with bidirectional data transfers. As shown in Figure 2 (left side), the KV cache in GPU memory cannot be freed or overwritten (step 2) until the swap to CPU memory (step 1) is complete, since the KV cache is updated after each token generation iteration. We empirically note that this dependency introduces frequent synchronization between CPU and GPU memory during KV cache swapping, leading to increased serving latency and reduced system throughput (see § 3.2).
Instead, we propose dynamic parameter remapping [^1] In this paper, swap refers to bidirectional data transfer to and from CPU/GPU memory, whereas remap denotes unidirectional data transfer from CPU to GPU memory., where we simply repurpose a portion of the GPU memory originally allocated for model parameters to support a larger KV cache, enabling the system to accommodate increased KV cache requirements. Unlike KV cache swapping, parameter remapping introduces negligible synchronization overhead, as model parameters remain constant during inference. As shown in Figure 2 (right), it is a non-blocking, unidirectional data transfer that does not require offloading parameters back to the CPU. This immutability simplifies resource management, as the transfer is both deterministic and fixed in size. Notably, parameter remapping is particularly well-suited for multi-tenant LLM serving: it enables efficient and proactive reuse of memory allocated to inactive models [^2] Prior work observed that models frequently enter idle periods without incoming requests. In production traces, over 60% of models exhibit more than 1,000 idle intervals exceeding 10 seconds each (Yu et al., 2025). We refer to models in such intervals as inactive models. by repurposing their parameter storage as KV cache for active models.
However, realizing a practical parameter remapping engine raises several open questions, including when to initiate and stop remapping, which models to target, how many layers to remap, and which specific layers to select. For example, suboptimal layer selection can trigger excessive CPU–GPU data transfers, resulting in a 12.7% reduction in throughput (§ 7.5). Failing to halt remapping in a timely manner can introduce unnecessary pipeline overhead (Hestness et al., 2015; Xu et al., 2024), negatively impacting system performance and increasing serving latency by up to 49% (§ 7.6). The complexity of these decisions is further amplified by the dynamic nature of system load and the variability in request sequence lengths across inputs (Yu et al., 2025; Stojkovic et al., 2024b; Patel et al., 2024). As a result, these decisions must be made adaptively at runtime, highlighting the need for a sophisticated memory management and data transfer engine within the LLM inference serving system (§ 3.4).
In this paper, we introduce MIRAGE, a Dynamic Remapping Engine that automatically and elastically adjusts KV cache size at layer [^3] In this paper, a layer refers to a hidden layer comprising feed-forward, attention, and normalization components, though the exact structure may vary across models (Agrawal et al., 2024a; Liakopoulos et al., 2025; Xu et al., 2024). For example, Llama-2-13B consists of 40 such hidden layers (Touvron et al., 2023b). granularity based on runtime demands. Such flexibility allows MIRAGE to be integrated into any LLM inference serving system with any schedulers (§ 8). We make the following key contributions:
  • We deeply study several potential GPU memory management optimization strategies to motivate dynamic parameter remapping.
  • We introduce MIRAGE, a Dynamic Remapping Engine for multi-tenant LLM inference serving systems that elastically adjusts the KV cache size at runtime by remapping parameter memory as KV cache memory.
  • We create a layer eviction algorithm that leverages the circular execution characteristics of LLMs to select which layers to remap. In addition to empirical evaluation, we provide a theoretical proof showing the optimality of our algorithm.
  • We integrate MIRAGE into vLLM to evaluate its effectiveness. Compared to the baseline vLLM, MIRAGE increases throughput by 7.9%-86.7% and reduces tail latency by 20.7%-99.3%.
Figure 2. Both KV cache swapping and parameter remapping expand the effective KV cache capacity by utilizing CPU memory as an extension of GPU memory. KV cache swapping incurs overhead due to frequent synchronization, as it requires backing up KV cache to CPU memory after each token generation once swap begins. Instead, parameter remapping is unidirectional with negligible synchronization overhead.
2. Background
This section provides background on the state-of-the-art inference serving system (§ 2.1), highlighting that KV cache is the primary bottleneck in both scenarios (§ 2.2).
2.1. LLM Inference Serving
Inference Execution Workflow: LLM inference serving consists of two phases: (a) the prefill phase, which processes the full input sequence in a single pass, and (b) the decode phase, which generates one output token at a time. Figure 3 illustrates the data placement and execution workflow for a typical LLM inference serving system (Kwon et al., 2023). The two primary contributors to GPU memory are LLM model parameters and the KV Cache (Pope et al., 2023; Kwon et al., 2023; Agrawal et al., 2024b; Prabhu et al., 2025). 1 Before the inference serving phase, the model parameters are transferred from CPU memory (DRAM) to GPU memory (HBM) (Li et al., 2023; Kwon et al., 2023; Agrawal et al., 2024b; Prabhu et al., 2025; Fu et al., 2024). After loading parameters, 2 input tokens are then loaded to GPU memory, where 3 tokens are transferred to GPU compute units (where LLM kernel executes) for processing during the prefill phase. During the decode phase, the KV cache is generated and expands with the sequence length. Throughout each iteration, the KV cache is transferred between 4 LLM kernels and 5 GPU memory. When the maximum sequence length is reached or an end-of-sequence (EOS) token is generated, the output tokens produced by the GPU compute units are first written to GPU memory 6, and then transferred back to CPU memory 7.
KV Cache: The KV cache stores intermediate results from previously generated tokens during the decode phase, eliminating the need to recompute previously generated tensor values. Its size scales linearly with both the batch size and request sequence length.
Performance Metrics: The performance of an LLM inference serving system is usually evaluated using two key latency metrics: TTFT (time-to-first-token) and TBT (time-between-tokens). TTFT measures the time taken to generate the first output token from the moment a request enters the system, reflecting the initial responsiveness of an inference serving system. TBT measures the interval between generating consecutive output tokens for a given request, which reflects the overall perceived smoothness of the response. Throughput is another critical metric to evaluate inference serving systems alongside latency, which quantifies the number of requests/tokens a system can handle within a specific time frame.
Figure 3. LLM inference serving on GPUs. GH200 features 450 GB/s CPU-GPU bandwidth, while H100 only has 64 GB/s. Swapping KV cache/parameter between CPUs and GPUs during runtime can fully utilize the increased bandwidth.
2.2. Memory Bottlenecks in LLM Serving
The memory demands of LLMs have increased at a much faster rate than GPU memory capacity, making memory a critical bottleneck in LLM inference serving. As model sizes grow, they require not only greater computational resources but also significantly more memory to accommodate both model parameters and KV cache (Kwon et al., 2023; Agrawal et al., 2024b). To address this challenge, recent works focused on efficiently managing KV cache, as its size grows linearly with the sequence length and batch size, which are unknown prior to runtime (Pope et al., 2023; Kwon et al., 2023; Lee et al., 2024a; Liu et al., 2024a; Prabhu et al., 2025; Jin et al., 2024; Adnan et al., 2024). For example, vLLM employs PagedAttention (Kwon et al., 2023), which leverages a paging mechanism to manage KV cache pages during runtime. However, when the KV cache exceeds the available GPU memory, recomputation becomes necessary. This scenario is difficult to anticipate, as the sequence length is unknown before runtime and request arrival patterns vary in production environments (Stojkovic et al., 2024a; Patel et al., 2024; Stojkovic et al., 2024b). This unpredictability also affects multi-tenant LLM inference systems. Regardless of the scheduling policy, systems using PagedAttention face the same challenge as single-model serving: runtime KV cache usage is inherently unpredictable. Recomputation can significantly degrade the performance of the affected sequence, especially by increasing the tail latency of the overall serving system. To mitigate this, a mechanism is desired to prevent recomputation when KV cache capacity is exceeded, thereby reducing tail latency in LLM inference serving.
3. Motivation
This section presents three approaches to expanding KV cache capacity, enabled by the significantly higher CPU–GPU bandwidth of modern hardware architectures like GH200.
  • Approach 1: offloading computation to CPUs (Jiang et al., 2024), where a portion of the attention operations is offloaded to CPUs, allowing the KV cache to be generated and stored in CPU memory instead of consuming GPU memory. As a result, the combined CPU–GPU system can accommodate larger batch sizes, leading to increased throughput (§ 3.1).
  • Approach 2: KV cache swapping (Xu et al., 2024), which swaps out portions of the KV cache to CPU memory when not immediately needed, freeing up GPU memory for active KV cache usage. This allows the system to accommodate larger batch sizes within the available GPU memory (§ 3.2).
  • Approach 3 (proposed): parameter remapping, which dynamically repurposes model parameter memory to automatically and elastically expand KV cache capacity at runtime, enabling support for larger batches or longer sequences (§ 3.3).
KV cache swapping (Approach 2) and parameter remapping (Approach 3) better utilize the available bandwidth without incurring substantial performance overhead and are more effective than offloading computation to CPUs (Approach 1), which is primarily suited to legacy systems with limited bandwidth. We further compare KV cache swapping (Approach 2) with parameter remapping (Approach 3) and show that parameter remapping has the potential to offer even better performance, particularly in multi-tenant LLM inference scenarios. We conclude with a quantitative analysis of computation and parameter loading times, highlighting challenges and the importance of efficient parameter remapping (§ 3.4).
3.1. Limitations of CPU Offloading
Prior works have demonstrated that offloading part of the attention computation and KV cache states from GPUs to CPUs can effectively increase the batch size and thus inference throughput in LLM inference serving systems (Jiang et al., 2024). Although offloading computation to CPUs has shown promising performance on earlier hardware platforms, we find it is less efficient on advanced systems such as the GH200 superchip, where limited CPU parallelism, rather than CPU-GPU bandwidth, is the bottleneck. Figure 5 presents the CPU computation time for processing OPT-13b across varying batch sizes serving requests from the ShareGPT dataset (Team, 2025). We compare two approaches: (a) offloading computations to CPUs, and (b) continuing GPU computation while remapping parameters to expand KV cache capacity to support large batch sizes. Parameter remapping is preferable when the parameter loading time does not exceed the CPU computation time. In GH200 systems, high CPU–GPU bandwidth but limited CPU parallelism make parameter loading significantly faster than CPU processing, reinforcing the suitability of parameter remapping. In contrast, on an H100, offloading computation to CPUs may be viable, even when compared to a low remapping percentage [^4] We define the remapping percentage as the fraction of memory used by model parameters that are remapped as added capacity for KV cache, while the remaining portion continues to be used for storing model parameters., such as 30%. Therefore, on systems with high CPU-GPU bandwidth, remapping parameters offers a more promising solution than offloading computation to CPUs.
(a) Temporal sharing (remap 2/3 param of inactive Model-B).
(b) Spatial sharing (remap 1/3 param of each model).
Figure 4. Remapping parameters for KV cache allows for larger batches when deploying two OPT-13b models on H100, in both (a) temporal sharing and (b) spatial sharing scenarios.
3.2. Limitations of Swapping KV Cache
With emerging hardware offering higher CPU–GPU bandwidth, KV cache for non-executing (future) layers can be temporarily swapped into CPU memory and brought back to GPU memory when needed. This approach enables larger batch sizes and enhances overall system performance, as demonstrated by Pie (Xu et al., 2024). However, KV cache swapping presents significant challenges, primarily because the KV cache must be updated after every token generation (Yu et al., 2022; Kwon et al., 2023; Agrawal et al., 2024b). This results in frequent synchronization between CPU and GPU memory, introducing overhead that affects serving latency. In addition to synchronization overhead, the bidirectional nature of data transfers can reduce effective CPU–GPU bandwidth, preventing the system from achieving an ideal pipeline of data transfer and computation. Our measurements of CPU host memory (LPDDR5X [^5] The trend observed in our measurements also holds for other memory technologies (e.g., DDR4, DDR5) beyond LPDDR5X (Esmaili-Dokht et al., 2024).) bandwidth show that unidirectional (read-only of CPU host memory) CPU-to-GPU transfers can reach $\sim$427 GB/s. However, when the read-to-write ratio is 1:1, bandwidth drops to $\sim$366 GB/s, resulting in a $\sim$15reduction. This bandwidth degradation can further impact serving throughput.
3.3. Our Proposal: Parameter Remapping
We introduce Parameter Remapping, where, instead of swapping KV cache, we propose to dynamically remap the space used for model parameters to automatically and elastically expand the KV cache size at runtime. Dynamic parameter remapping reclaims GPU memory by repurposing a portion of the model parameters for KV cache usage during runtime, enabling online inference serving systems to handle peak loads when GPU memory becomes insufficient for KV cache. As shown in the two examples in Figure 4, deploying two OPT-13b models on an H100 GPU using either temporal (Figure 4) or spatial (Figure 4) sharing one GPU consumes approximately 52GB of the available 80GB memory for model parameters (Kwon et al., 2023), leaving only about 28GB for KV cache. By remapping one-third of the model parameters as KV Cache—either $2/3$ from the inactive model in temporal sharing or $1/3$ from each model in spatial sharing—the maximum batch size can be increased by up to 1.75×1.75$\times$1.75 ×, resulting in a corresponding 1.75×1.75$\times$1.75 × improvement in system throughput. Unlike KV cache swapping, parameter remapping involves non-blocking, unidirectional data transfers and does not require backing up model parameters from GPU to CPU, as parameters remain unchanged during inference. As a result, synchronization overhead is significantly reduced. Additionally, parameter remapping is particularly effective in multi-tenant LLM inference scenarios, such as multi-agent workflows (Talebirad and Nadiri, 2023; Chan et al., 2023; Li et al., 2024b; Han et al., 2024b; Kim et al., 2025) and production deployments where many models remain idle for long periods of time (Yu et al., 2025). In such settings, memory allocated to inactive models can be reclaimed for active ones, a flexibility that KV cache swapping lacks, as inactive models may have no KV cache to be swapped.
(a) CPU Compute.
(b) GPU Compute.
Figure 5. (a) Computation time of different batch sizes when offloading to CPUs. Parameter remapping becomes a more promising approach when CPU-GPU bandwidth is sufficient. (b) GPU computation time fluctuates with batch size based on the load in an online serving system, requiring the parameter remapping percentage to be adjusted dynamically.
3.4. Challenges in Parameter Remapping
While parameter remapping can improve performance, it introduces potential overhead and requires careful design. To avoid impacting inference latency and throughput, parameter loading must ideally overlap with GPU computation, ensuring that the parameters of each layer are fully loaded before its execution begins. For example, while the GPU is executing layer 1, the parameters for layer 2 can be loaded concurrently from CPU to GPU memory. The ideal case is when layer 2 finishes loading before layer 1 finishes execution, allowing GPU memory to hold only two layers at a time (Sheng et al., 2023; Rasley et al., 2020), unlike approaches that preload the entire model into GPU memory. Although GH200 offers significantly higher CPU–GPU bandwidth than previous architectures, it remains insufficient for full on-demand remapping, as the execution of a layer is faster than transferring its parameters from CPU to GPU memory. As a result, only a fraction of the parameters—defined by the remapping percentage—can be dynamically transferred at runtime, while the rest must be preloaded into GPU memory before the inference serving starts. In this section, we examine the challenges involved in making these design decisions.
When to trigger remapping? While an ideal pipeline can fully overlap GPU computation with parameter loading from CPU to GPU memory, the parameter remapping engine needs to initiate parameter remapping only when necessary, as even a perfectly pipelined approach introduces inherent overhead (Hestness et al., 2015). Likewise, it is important to halt parameter remapping once it is no longer needed. Delaying this decision can increase the latency of the serving system by up to 49% (§ 7.6). These decisions require dedicated mechanisms and careful design for online monitoring of the LLM inference serving system.
Which models to remap? In multi-tenant LLM inference serving systems, not all models are active simultaneously (Patke et al., 2024; Yu et al., 2025), which presents an opportunity for the parameter remapping engine to prioritize remapping parameters of inactive models. However, this approach requires careful design, as overly aggressive offloading may lead to starvation—especially given the unpredictable nature of request arrival patterns in online serving environments (Stojkovic et al., 2024b). An inefficient design is likely to prevent certain models from responding promptly; for example, a suboptimal model selection can lead to a 22.0% increase in serving tail latency (§ 7.6).
How many layers to remap? Determining the optimal remapping percentage is challenging due to the dynamic nature of request arrival rates in real-world online inference serving (Stojkovic et al., 2024a; Patel et al., 2024; Stojkovic et al., 2024b). The fluctuations in request arrival rates can affect the GPU computation time, which in turn affects the percentage of parameters that can be transferred without performance degradation. We illustrate the need for dynamic adjustment of the remapping percentage using an example of serving OPT-13b with varying batch sizes on the ShareGPT dataset (Team, 2025). Figure 5 shows how the GPU computation time for generating one token in the decode phase varies with batch size, while parameter load time—determined solely by the remapping percentage—remains constant. The intersection points of these curves represent configurations where parameter loading is fully overlapped with computation, maximizing efficiency. As shown, the optimal remapping percentage varies with changes in the runtime batch size, reflecting shifts in system load. Using a static remapping percentage may result in suboptimal resource use: underutilizing GPU memory at off-peak load and causing compute stalls at peak load. For example, a 30% remapping percentage tuned for average batch size in ShareGPT results in CPU–GPU bandwidth becoming a bottleneck during off-peak workloads. Therefore, the remapping percentage must be dynamically adapted at runtime.
Which layers to target? Once the remapping percentage is set, the next design challenge is determining which specific layers to remap. This remains non-trivial, as MIRAGE must decide how to allocate limited GPU memory across all layers—determining which layers remain permanently in GPU memory and which are transferred on the fly. To achieve an efficient pipeline, a layer selection strategy is needed to guide the layer-sharing mechanism in MIRAGE. A suboptimal layer selection can lead to a reduction in serving throughput by 12.7% (§ 7.5).
4. MIRAGE Overview
We introduce MIRAGE, a Dynamic Remapping Engine for multi-tenant LLM inference serving systems that elastically adjusts the KV cache size by remapping parameter memory as KV cache memory. Figure 6 shows the MIRAGE architecture, and depicts how MIRAGE can be integrated into existing LLM inference serving systems.
4.1. Architecture of MIRAGE
MIRAGE is designed to have three key components, marked in green color in Figure 6.
Metadata Store: The Metadata Store maintains the information about the models and their memory utilization. Model information includes the currently scheduled models (referred to as active models) and idle models currently not receiving requests (referred to as inactive models). Memory utilization information captures KV cache usage, with a focus on the amount of available free memory.
Remapping Controller: The Remapping Controller dynamically repurposes GPU memory allocated to inactive model parameters for use as KV cache by active models. When KV cache demand exceeds available GPU memory, the Remapping Controller reclaims a portion of the parameter memory to expand KV cache capacity. Conversely, when KV cache usage subsides, the reclaimed memory is remapped back for parameter storage.
Async Transfer Engine: The Async Transfer Engine is responsible for loading parameters and coordinating synchronization between GPU computation and CPU-GPU data transfer. When the Remapping Controller initiates parameter remapping, a portion of GPU memory originally allocated for model parameters is repurposed for KV cache usage. If the reclaimed parameters are later required for computation, the Async Transfer Engine fetches them back from CPU memory to GPU memory. By leveraging the deterministic execution order of layers within each decode iteration, the Async Transfer Engine overlaps CPU–GPU data transfers with ongoing GPU computation.
Figure 6. Overview of MIRAGE.
4.2. Workflow
Figure 6 depicts the workflow of the interaction between various components to achieve the goal of dynamic parameter remapping in MIRAGE. The Scheduler [^6] MIRAGE is compatible with a wide range of popular scheduling policies (e.g., either spatial or temporal sharing of GPU resources for multi-tenant LLM inference serving). The scheduling policy design itself is beyond the scope of this paper. selects which models to serve, referred to as active models. Once active models are chosen, the Memory Allocator updates the parameter and KV cache memory usage accordingly. Using both model information and KV cache utilization data, the Remapping Controller determines whether parameter memory needs to be remapped to expand KV cache capacity. If remapping is required, the Remapping Controller updates memory usage records in the Memory Allocator and triggers the Async Transfer Engine to initiate data transfer. The GPU then executes the LLM inference tasks. With support from MIRAGE, a portion of parameter memory in GPUs can be dynamically reclaimed at runtime, allowing the system to accommodate larger batches or longer sequences.
5. Remapping Controller
Figure 7. Remapping Controller decides which model and layer to remap based on runtime KV cache usage (Model-A is active).
MIRAGE enables dynamic, runtime management of GPU memory through its Remapping Controller. Figure 7 illustrates an example execution workflow of the Remapping Controller across five phases, each representing a snapshot of model activity within an LLM serving system. Depending on runtime conditions, the Remapping Controller decides either to initiate remapping, continue remapping additional layers from the active or inactive models, or halt remapping altogether (reclaimed KV cache memory restored for parameter storage). In this section, we describe how Remapping Controller answers the key questions, including when to remap (§ 5.1), which models to remap (§ 5.2), how many (§ 5.3), and which layers to remap (§ 5.4). Putting it all together, we then present an algorithm that can seamlessly integrate MIRAGE into existing LLM inference serving frameworks, enabling dynamic remapping of model parameter memory to automatically and elastically expand KV cache capacity at runtime (§ 5.5).
5.1. When to Remap?
The Remapping Controller initiates parameter remapping when the KV cache capacity is exhausted. While unidirectional parameter remapping incurs lower overhead than bidirectional KV cache swapping, MIRAGE activates it only when necessary. As such, accurately determining both the initiation and termination points of remapping is critical. Given the dynamic nature of real-world production workloads (Cortez et al., 2017; Hadary et al., 2020; Shahrad et al., 2020; Stojkovic et al., 2024a; Patel et al., 2024; Stojkovic et al., 2024b), the Remapping Controller enables remapping during peak request periods to accommodate larger batch sizes. During non-peak periods, when KV cache space is sufficient, remapping is disabled to avoid the overhead of reloading previously evicted parameters into GPU memory. The reclaimed KV cache memory is then restored for parameter usage.
5.2. Which Models to Remap?
This section explains how the Remapping Controller determines which models to remap. We begin by reviewing existing GPU sharing techniques used in multi-tenant inference serving systems, as the Remapping Controller adapts its model selection strategy based on the specific sharing mechanism employed. We then present a detailed explanation of our model selection strategy.
GPU sharing background: Given the memory capacity of emerging GPUs, enabling GPU sharing is essential to fully utilize their processing capabilities. Prior works have examined both temporal (Bai et al., 2020; Gujarati et al., 2020; Xiao et al., 2020; Fu et al., 2024; Patke et al., 2024) (example in Figure 4) and spatial (Dhakal et al., 2020; Romero et al., 2021; Lee et al., 2024b; Strati et al., 2024; Choi et al., 2022; Xu et al., 2022; Han et al., 2024a; Wang et al., 2024a) (example in Figure 4) sharing strategies for efficient machine learning inference serving on GPUs. Temporal sharing allocates the entire GPU to one model at a time, allowing for larger batches per model. This approach is especially useful in multi-agent workflows—where the output of one model becomes the input to another (Talebirad and Nadiri, 2023; Chan et al., 2023; Li et al., 2024b; Han et al., 2024b; Kim et al., 2025)—and in cloud production environments, where models often experience idle periods with no incoming requests (Yu et al., 2025). In contrast, spatial sharing divides GPU resources among multiple models simultaneously, allowing for faster response times at the expense of smaller batch sizes per model (Lee et al., 2024b; Duan et al., 2024). Each GPU resource sharing strategy is tailored to specific use cases and can be selectively employed by multi-tenant LLM inference schedulers. The Remapping Controller in MIRAGE is designed to be compatible with any scheduling policy, although the design of the scheduler itself is beyond the scope of this study.
Model selection strategy: We now present the model selection strategy to determine which models to remap parameter layers from. The Remapping Controller prioritizes remapping parameters from inactive models with the lowest priority, as determined by the scheduling policy, rather than from active ones. Under temporal GPU resource sharing, models scheduled for future execution are considered inactive. In contrast, under spatial sharing, all models are treated as active, requiring each to remap parameters. The number of layers to remap and their selection are determined using the strategies described in § 5.3 and § 5.4. If the scheduler does not specify model priorities, the Remapping Controller defaults to a round-robin policy, treating all models equally. In this case, it applies a Most Recently Used (MRU) strategy, remapping parameters from the most recently activated model to defer parameter transfer costs as far into the future as possible. This strategy is especially effective when future request patterns are unpredictable or when model priorities are unspecified.
Example: Phase 2 of Figure 7 illustrates an example of the model selection policy. In this scenario, Model-A is active, while Model-B and Model-C are inactive, with the scheduler assigning the lowest priority to Model-C. The Remapping Controller first evicts and remaps parameters from Model-C. To avoid starving inactive models, a maximum remapping threshold is enforced (see § 5.3), ensuring that enough parameters remain in GPU memory to support cold-start loading during the prefill phase when a model becomes active. After reaching the remapping limit for Model-C, the Remapping Controller continues with Model-B. Only once all inactive models have reached their remapping thresholds does the Remapping Controller begin remapping parameters from the active Model-A.
5.3. How Many Layers to Remap?
To hide parameter transfer latency by overlapping it with GPU computation, the transfer from CPU to GPU must finish before the computation completes; otherwise, it becomes a performance bottleneck. Since both model parameter sizes and CPU–GPU bandwidth are known ahead of runtime, the transfer time per layer can be profiled offline and treated as prior knowledge by the online inference system. As a result, the Remapping Controller only needs to monitor GPU computation time, denoted as $T_{Compute}$, at runtime to determine whether additional parameters can be remapped. For an active model, $T_{Compute}$ refers to the time required to generate the current token during the decode phase. For an inactive model, it represents the duration of the prefill phase, which is more compute-intensive and typically longer, as no tokens have been generated yet (Patel et al., 2024; Agrawal et al., 2024b). The longer $T_{Compute}$ in the prefill phase allows more layers in inactive models to be remapped compared to when the model is active. Let $T_{T}$ denote the time required to transfer a single layer of parameters. To prevent parameter loading from becoming a new bottleneck, the number of remapped layers $N$ must satisfy the constraint $T_{T}\cdot N\leq T_{Compute}$.
5.4. Which Layers to Target?
This section details the layer selection strategy in the Remapping Controller. We begin by describing the strategy itself, followed by an example to illustrate its behavior. Finally, we provide a theoretical proof to demonstrate that the strategy we used is optimal.
Layer selection strategy: Understanding the layer selection strategy requires recognizing the circular execution pattern of layers in LLM inference, as shown in Figure 7. The auto-regressive nature of LLM inference makes each of the subsequent tokens execute the same layers again. Thus, LLM inference can be viewed as a circular execution of layers, where the execution of the last layer for the current token is succeeded by the execution of the first layer for the next token. The Remapping Controller uses a uniform-interval layer selection strategy, in which evenly spaced layers are chosen to share a common GPU memory region allocated for their parameters. These layers alternate access to the shared region—at any given moment, some layer parameters reside in GPU memory, while others remain in CPU memory. When a layer not currently in GPU memory is needed, its parameters are loaded into the shared region, replacing those of the previously loaded layer.
Example: To provide a more intuitive explanation, we use the example of Model-C with 8 layers (Phase 2 of Figure 7). The Remapping Controller employs a uniform-interval layer selection strategy. For instance, when remapping the parameters of a single layer, Layers 1 and 5 are selected to share a common GPU memory region (the other 6 layers remain in GPU memory permanently), alternating between CPU and GPU memory. The uniform-interval layer selection strategy is designed to balance serving latency across successive token generations. In contrast, ignoring the circular nature of LLM inference—by treating each token generation as an independent execution—may lead to suboptimal choices, such as selecting Layer 1 and Layer 8. Although this configuration allows sufficient time to load the parameters of Layer 8 after executing Layer 1, it does not provide enough time to reload the parameters of Layer 1 before it is needed again. This demonstrates the benefit of uniform-interval selection in circular execution, which ensures that each remapped layer can be restored to GPU memory in time for reuse.
Theoretical proof: Consider an n-layer model in which the Remapping Controller remaps the parameter memory of $\alpha$ layers for KV cache. Consequently, all $n$ layers must be scheduled within the GPU memory footprint of only $n-\alpha$ layers. Suppose $m$ layers are transferred between CPU and GPU during each token generation. We note $L_{1},L_{2},\ldots,L_{m}$ as the $m$ layers (remapped layers in Figure 7) that are transferred from CPU to GPU. We define $T_{T}$ as the time taken to load parameters of a single layer from CPU to GPU memory, and $T_{c}$ as the time spent on computation for a single layer on GPU [^7] LLM architectures often adopt layers of uniform size, both in terms of parameters and compute requirements (Agrawal et al., 2024a; Liakopoulos et al., 2025), resulting in uniform parameter transfer and computation latency across layers. Therefore, $T_{C}$ and $T_{T}$ are equal for all (hidden) layers.. Our layer selection policy aims to satisfy two key objectives: (1) ensure that parameter transfer latency does not become a performance bottleneck; and (2) minimize CPU-GPU data transfer traffic.
To satisfy objective (1): the CPU-to-GPU parameter loading must be fully overlapped with GPU computation. During each token generation, the Remapping Controller transfers $m$ layers from CPU to GPU memory. The goal is to select $m$ layers such that the computation between their executions is sufficient to hide the associated transfer latency. For example, after layer $L_{1}$ completes execution, the Remapping Controller begins transferring parameters of layer $L_{2}$. The cumulative computation time of the layers executed between $L_{1}$ and $L_{2}$ must be long enough to mask the transfer time of $L_{2}$. Similarly, each subsequent transfer must also meet the condition to hide its transfer time, continuing until the parameters for layer $L_{1}$ (for the next token) are fully transferred. We formally present this constraint as:
(1) $T_{T}\leq\sum_{i=L_{1}}^{L_{2}}T_{c}\;\cdots,\;T_{T}\leq\sum_{i=L_{m-1}}^{L_{m}}T_{c},\ T_{T}\leq\sum_{i=L_{m}}^{L_{1}}T_{c}$
Consider that there are $k_{1}$ layers between $L_{1}$ and $L_{2}$, and $k_{m}$ layers between $L_{m}$ and $L_{1}$ (for the next token), we have:
(2) $\sum_{i=L_{1}}^{L_{2}}T_{c}=k_{1}\times T_{c},\;\cdots\;,\sum_{i=L_{m}}^{L_{1}}T_{c}=k_{m}\times T_{c}$
Leveraging Equations 1 and 2, we conclude that:
(3) $T_{T}\leq\min(k_{1}\times T_{c},\;\cdots,\;k_{m-1}\times T_{c},\;k_{m}\times T_{c})$
The Right-Hand Side (RHS) of Equation 3 denotes the threshold beyond which transfer time becomes a performance bottleneck. To maximize the RHS under the constraint $k_{1}+k_{2}+\;\cdots\;+k_{m}=n$, the optimal solution is achieved when $k_{1}=k_{2}=\cdots=k_{m}$. In other words, within the circular layer execution model, the selected $m$ layers must be evenly spaced, ensuring uniform intervals between consecutive remapped layers. This aligns with the geometric principle that, for $n$ points on a circle, the configuration maximizing the minimum pairwise distance is achieved by placing them at equal angular intervals.
To satisfy objective (2): we aim to find the smallest valid integer value of $m$. We begin with $m=\alpha$, meaning that only $\alpha$ layers are loaded from CPU memory during each token generation. However, this configuration is infeasible because loading a remapped layer into GPU memory requires evicting a resident layer, which will be needed in the next token generation due to the circular reuse pattern of all layers. As a result, $m=\alpha$ is insufficient, and the minimum viable value becomes $m=\alpha+\beta$ where $\beta\geq 1$. In this configuration, $n-(\alpha+\beta)$ layers remain permanently in the GPU memory, while the remaining $\alpha+\beta$ layers share a GPU memory region sized for $\beta$ layers.
We start with $\beta=1$. To satisfy design objective (1), we have Equation 3 to make the data transfer time shorter than the computation time. However, the prerequisite of Equation 3 is that parameter transfers from CPU to GPU can be initiated at any time, overlooking the data dependencies between layers: the $\alpha+1$ layers share a single GPU memory slot, meaning that the transfer of a layer cannot begin until its computation has completed. Accounting for this constraint, the time budget becomes more restrictive:
(4) $T_{T}\times(\alpha+1)\leq T_{c}\times(n-\alpha-1)$
To relax this constraint, one possible solution is to adopt a double-buffering approach, increasing the number of layers transferred on the fly from $\alpha+1$ to $\alpha+2$. With two GPU memory slots, these $\alpha+2$ layers can be transferred and computed in a staggered manner, allowing the transfer of the next layer to begin without waiting for the execution of the current layer to complete. This results in a less restrictive timing condition:
(5) $T_{T}\times(\alpha+2)\leq T_{c}\times n$
When $\alpha$ is small and $n$ is large, setting $m=\alpha+1$ is generally preferred as it minimizes CPU-GPU transfer. In contrast, $m=\alpha+2$ is more suitable for larger models with more remapped parameters. For instance, based on Equations 4 and  5, when $\alpha\geq 9$ for a model with 40 layers ($n=40$), choosing $m=\alpha+2$ can provide better performance (§ 7.5).
5.5. Putting All Together
Algorithm 1 outlines the complete execution workflow of the Remapping Controller. Consistent with existing LLM inference serving schedulers, the Remapping Controller makes parameter remapping decisions at a per-token granularity. At each iteration step, it first evaluates the runtime status of the Memory Allocator, to determine whether remapping needs to be triggered (Line 3). Then, it marks the GPU memory previously occupied by parameters as available for KV cache. The Remapping Controller then updates the list of model layers whose parameters will be remapped between CPU and GPU memory (Line 4). As illustrated in the remapping() function of Algorithm 1, the Remapping Controller continuously tracks the current remapping percentage for each model (Lines 18–23) and prioritizes remapping parameters from the model (Line 16) with the lowest priority decided by the scheduler. Meanwhile, when the KV cache space is freed after requests finish, the Remapping Controller updates the remapping percentage based on the available memory information from the memory allocator (Lines 7-12). Finally, the Remapping Controller updates parameter remapping information (remapped models and layers) to the GPU LLM Kernel, which executes the current iteration accordingly (Line 13).
1 Function Remapping_Controller():
2       foreach step() $\in$ LLM Inference Serving Scheduler do
3             if Running out of KV Cache blocks then
4                   remapped_layer_list = remapping();
5                   enable_remap = True;
6
7             else
8                   foreach Tensor $\in$ Added KV Cache do
9                         if Tensor is empty then
10                               remapped_layer_list = remapping();
11
12
13                   end foreach
14                  if All Tensors are empty then
15                         enable_remap = False;
16
17
18             GPU LLM Kernel(enable_remap, remapped_layer_list);
19
20       end foreach
21
22
23
24Function Remapping():
25       model = pop(inactive_model_list);
26       if model is not NULL then
27             model.remapped_layers++;
28             if model.remapped_layers == model.layers then
29                   inactive_model_list.remove(model);
30
31             sort(inactive_model_list);
32
33       else
34             remapping parameters of active models;
35
36
37
Algorithm 1 Remapping Controller Workflow
6. Implementation
We integrate MIRAGE into vLLM (Kwon et al., 2023), a state-of-the-art LLM inference-serving framework. We add $\sim$3,00030003,0003 , 000 lines of Python code and $\sim$300300300300 lines of C++/CUDA code to enable support for parameter remapping engines and dynamic KV cache allocation.
Remapping Controller and Async Transfer Engine: During inference, the Remapping Controller coordinates with the Transfer Engine to specify which layers from which models are to be remapped. If remap is disabled, MIRAGE follows the standard execution flow without modification. When remap is enabled, the Transfer Engine initiates an asynchronous transfer of the parameters of the target layer, overwriting [^8] State-of-the-art LLM inference frameworks, such as vllm (Kwon et al., 2023), keep a complete copy of all model parameters in CPU memory before transferring to GPU memory. the GPU tensor memory previously used by layers that have completed execution. To ensure correctness, the GPU performs memory barrier synchronization checks before launching kernels that depend on remapped parameters.
KV Cache Allocation: Most state-of-the-art LLM frameworks, such as vLLM (Kwon et al., 2023), support dynamic management of KV cache memory space. However, the physical memory for KV cache still needs to be statically allocated before serving starts. However, during parameter remapping, a portion of the parameter memory must be dynamically reconfigured as KV cache memory, requiring support for runtime dynamic KV cache allocation and deallocation. To enable dynamic allocation and remapping of GPU memory, we adopt the vAttention design (Prabhu et al., 2025), which reserves sufficient virtual memory space for the KV cache and maps physical memory pages on demand. Once parameter tensors are released, the KV cache engine can immediately reuse the freed physical memory. This approach also allows GPU physical KV cache memory to be efficiently shared across multiple models.
7. Evaluation
We comprehensively evaluate the effectiveness of MIRAGE by answering the questions below:
  • How effective is MIRAGE in reducing tail latency and improving throughput under temporal GPU sharing? (§ 7.2)
  • Is MIRAGE still effective in enhancing performance under spatial GPU resource sharing? (§ 7.3)
  • How effective is the parameter remapping mechanism compared with KV cache swapping? (§ 7.4)
  • How effective is the layer selection strategy used? (§ 7.5)
  • How important is the decision of when to remap and how many layers to remap? (§ 7.6)
Some highlights of our results include:
  • MIRAGE significantly improves multi-tenant LLM inference with temporal GPU sharing, reducing tail TBT latency by 44.8%–82.5%, tail TTFT latency by 74.8%–99.3%, and achieving 39.9%–86.7% higher throughput compared to vLLM.
  • In addition to temporal sharing of GPU resources, MIRAGE also supports spatial sharing. MIRAGE achieves 20.7%-65.5% reduction in tail latency.
  • Compared to KV cache swapping, MIRAGE achieves 47.1% higher throughput. Additionally, it supports multi-tenant LLM inference serving, providing greater flexibility.
Table 1. Evaluated model combinations and their corresponding GPU memory reservations (of total GPU memory).
C1 OPT-13b (35%), Lllma-2-13b (35%), Lllma-3-8b (20%)
C2 OPT-30b (65%), OPT-6.7b (15%)
(a) MIRAGE reduces tail TBT latency.
(b) MIRAGE reduces tail TTFT latency.
(c) MIRAGE improves throughput.
Figure 8. MIRAGE reduces latency and improves throughput compared with vllm on both C1 and C2.
7.1. Experimental Setup
Platforms: We integrate MIRAGE into vLLM v0.7.3 (Kwon et al., 2023), using CUDA 12.4, PyTorch 2.5.1, and Linux kernel 5.14. All experiments were conducted on the NVIDIA GH200 Grace Hopper Superchip, which features an H200 GPU with 96GB of HBM3 memory and a CPU host with 72 Arm Neoverse V2 cores and 224GB of LPDDR5X memory, connected via a 900 GB/s NVLink interconnect (Nvidia, 2025c).
Models: We evaluated the following models: OPT (Zhang et al., 2022) with 13b and 30b parameters; Llama-2 (Touvron et al., 2023a) with 13b parameters; and Llama-3 (Dubey et al., 2024) with 8b parameters. For multi-tenant serving, we evaluate two model combinations C1 and C2 in Table 1.
Traces: We evaluate MIRAGE using the Azure coding LLM inference traces that capture real-world query arrival patterns characterized by bursty query patterns (Patel et al., 2024; Stojkovic et al., 2024b). Following prior work (Ahmad et al., 2024; Gujarati et al., 2020; Li et al., 2023; Miao et al., 2024), we scale these traces to different query arrival rates while preserving their original characteristics, such as burstiness.
Datasets: We use the ShareGPT (Team, 2025) and Alpaca (Taori et al., 2023) datasets, both of which contain input-output pairs from real LLM services. To simulate realistic multi-tenant scenarios, we additionally generate requests of varying lengths (synthetic datasets).
Baselines: We compare MIRAGE against two baselines: vLLM (Kwon et al., 2023) and Pie (Xu et al., 2024). vLLM is a state-of-the-art LLM inference serving system widely used in both academia and industry. Pie is the first to explore using the NVIDIA Grace Superchip for KV cache swapping.
Metrics: Our evaluation focuses on two key metrics: (a) tail (P99) latency, which is critical for managing congestion in online systems (Haque et al., 2015; Prekas et al., 2017; Delimitrou and Kozyrakis, 2018; Kaffes et al., 2019; Zhao et al., 2023); we consider both TBT and TTFT latency; and (b) throughput (tokens per second), which quantifies the number of tokens a system can handle within a specific time frame.
7.2. MIRAGE Supports Temporal GPU Sharing
We first evaluate MIRAGE under temporal GPU resource sharing, which is well-suited for multi-agent workflows and scenarios with frequent model idle periods.
Tail Latency: We begin by evaluating the effectiveness of MIRAGE through a comparison of tail TBT and TTFT latencies. Figures 8(a) and 8(b) present the P99 TBT and TTFT latency comparisons between MIRAGE and vLLM. Across all scenarios, MIRAGE consistently reduces tail latency. On average, when serving C1, MIRAGE achieves an average reduction of 54.4% in P99 TBT latency and 96.7% in P99 TTFT latency across both the Alpaca and ShareGPT datasets. Similarly, when serving C2, MIRAGE achieves an average reduction of 44.8% in P99 TBT latency and 74.8% in P99 TTFT latency on the same datasets.
Without MIRAGE, peak-load conditions that exhaust available KV cache typically cause existing serving systems (Kwon et al., 2023) to pause the decode process for all active requests, evict one, and recompute its KV cache to free space. MIRAGE eliminates the need for recomputation, maintaining low tail latency, thereby ensuring stable performance for online inference serving.
Figure 9. MIRAGE outperforms vLLM on OPT-30b (Model-A) + OPT-6.7b (Model-B) with different request arrival rates.
Throughput: By using the Async Transfer Engine, MIRAGE overlaps parameter transfer with GPU computation. This not only reduces tail latency but also sustains high serving throughput. Across all evaluated scenarios shown in Figure 8(c), MIRAGE consistently increases throughput. On average, when serving C1, it achieves a 39.9% improvement in throughput across both the Alpaca and ShareGPT datasets. Similarly, when serving C2, MIRAGE delivers an average throughput gain of 45.3% on the above datasets.
Figure 10. MIRAGE outperforms vLLM on C2 across both short (S) and long (L) input requests.
Varied Arrival Rates: In real-world multi-tenant scenarios, different models or tasks often experience varying arrival rates (Romero et al., 2021; Li et al., 2023; Sun et al., 2024; Yu et al., 2025). MIRAGE is designed to handle this variability effectively. To demonstrate its effectiveness, we evaluate MIRAGE by serving C2, with different arrival rates across different models. As shown in Figure 9, MIRAGE consistently improves performance across all tested scenarios of varied input arrival rates, achieving an average of 86.7% increase in throughput and 82.5% and 99.3% reduction in TBT and TTFT tail latencies, respectively.
Varied Inputs: To emulate realistic multi-tenant scenarios—where different models handle distinct tasks within a multi-agent flow and often operate on separate input datasets (Li et al., 2024b, a; Talebirad and Nadiri, 2023; Sun et al., 2024)—we construct arrival traces for two models (evaluated on C2) using two request-type combinations: (a) synthetic long requests (average of 1734 tokens) paired with synthetic short requests (average of 634 tokens); and (b) synthetic short requests paired with synthetic long requests. As shown in Figure 10, MIRAGE consistently improves performance across all evaluated scenarios, achieving an average throughput increase of 65.6% and reductions of 57.1% and 98.1% in P99 TBT and TTFT latencies, respectively.
Which models to remap? We evaluate the effectiveness of our model selection strategy in Remapping Controller (§ 5.2) under temporal GPU sharing. This strategy aligns with the multi-tenant scheduling policy, which determines the execution order of models. In the absence of explicit model priorities, MIRAGE leverages a default round-robin scheduling policy combined with a Most Recently Used (MRU) eviction strategy to maintain performance. To assess its effectiveness, we compare MRU against a Least Recently Used (LRU) policy, which remaps the parameters of the model most likely to be served next. Figure 11 presents performance results for the C1 model combination under round-robin execution. With the MRU policy, MIRAGE achieves up to a 22.0% reduction in tail latency compared to the LRU policy. MRU consistently outperforms LRU by deferring transfer costs, remapping parameters of models that are less likely to be reused in the near future.
Figure 11. MRU (used in MIRAGE) remaps models expected to be used furthest in the future, reducing tail latency and improving throughput compared to LRU.
7.3. MIRAGE Supports Spatial GPU Sharing
In addition to supporting temporal GPU sharing, MIRAGE also performs effectively under spatial sharing for online inference. In this setup, all models remain active and maintain separate KV cache spaces. To assess the effectiveness of MIRAGE, we evaluate the model combination C1 on both Alpaca and ShareGPT datasets using Azure coding traces. We evaluate the performance of MIRAGE under two representative spatial sharing methods on NVIDIA-based GPU systems: (1) Non-strict physical isolation, as seen in Multi-Process Service (MPS) (Nvidia, 2025f), enables concurrent execution by allowing multiple processes to share internal GPU resources. It offers limited isolation since it does not partition components such as caches or memory controllers. (2) Strict physical isolation, exemplified by Multi-Instance GPU (MIG) (Nvidia, 2025e), partitions the GPU into independent instances, providing strong physical isolation across workloads.
Figure 12. MIRAGE achieves higher throughput and lower tail TBT latency than vLLM under non-strict physical isolation.
Non-Strict Physical Isolation: As shown in Figure 12, MIRAGE effectively reduces both tail TBT and TTFT latency as request arrival rates increase toward peak load. On average, it achieves a 65.5% reduction in P99 TBT, a 20.7% reduction in P99 TTFT, and a 6.6% increase in throughput across varying arrival rates. These results highlight the robustness of MIRAGE: it not only delivers strong performance under temporal GPU resource sharing but also offers significant improvements in spatial sharing configurations.
Strict Physical Isolation: MIRAGE also maintains strong performance under strict physical isolation. As shown in Figure 13, it significantly reduces both tail TBT and TTFT latency as arrival rates increase, achieving an average 57.4% reduction in P99 TBT, 34.8% reduction in P99 TTFT, and a 7.9% increase in throughput across different models and load conditions. Notably, strict physical isolation between models can be viewed as a special case of single-model LLM inference. This further demonstrates the flexibility and robustness of MIRAGE across diverse deployment scenarios.
(a) Tail TBT Latency.
(b) Tail TTFT Latency.
(c) Throughput.
Figure 13. Comparison between vllm and MIRAGE under non-strict physical isolation. MIRAGE achieves slightly higher throughput, but reduces tail TBT and TTFT latency significantly.
Figure 14. While MIRAGE and Pie exhibit similar tail TBT latency, MIRAGE achieves lower tail TTFT latency and delivers higher throughput.
7.4. MIRAGE vs KV Cache Swapping
Compared to KV cache swapping, MIRAGE achieves higher CPU–GPU bandwidth through unidirectional data transfer (§ 3.2), resulting in increased throughput. We evaluate MIRAGE against Pie (Xu et al., 2024), a KV cache swapping based method. Figure 14 shows the tail latency and throughput for serving the OPT-13b model on the Alpaca dataset of vLLM, Pie, and MIRAGE. The GPU memory utilization for single-model serving follows the values listed in Table 1, which also apply to § 7.5 and § 7.6. MIRAGE matches Pie in tail TBT latency, while reducing tail TTFT latency by 93.6% and improving throughput by 47.1%. This improvement is attributed to the ability of MIRAGE to avoid the overhead of backing up KV cache to CPU memory, thereby enabling more efficient utilization of CPU-GPU bandwidth.
Beyond performance improvements, MIRAGE provides greater flexibility in temporal sharing scenarios compared to KV cache swapping based approaches. In these scenarios, KV cache may be maintained only for models actively generating output tokens, while inactive models may not retain any. As a result, KV cache swapping based methodologies cannot reclaim memory from inactive models, limiting the ability to dynamically expand GPU memory for the KV cache of active models. MIRAGE overcomes this limitation by repurposing the memory allocated to inactive models, even in the absence of KV cache.
Figure 15. Using $m=\alpha+2$ (B/C) yields lower latency and higher throughput than consistently using $m=\alpha+1$ (A).
7.5. Effectiveness of Our Layer Selection Strategy
The layer selection strategy is important for enabling MIRAGE to effectively pipeline GPU computation with CPU–GPU parameter transfer, thereby minimizing performance overhead. The main trade-off involves selecting the number of layers ($m$, as defined in § 5.4) to transfer from CPU to GPU when $\alpha$ layers are remapped. Using $m=\alpha+1$ reduces the amount of data transfer, while $m=\alpha+2$ better addresses data dependencies by allowing full pipelining through a double-buffering mechanism. We present an empirical evaluation of three configurations: (A) using $m=\alpha+1$ consistently, (B) using $m=\alpha+2$ consistently, and (C) a dynamic scheme that selects $m=\alpha+1$ for small $\alpha$ and switches to $m=\alpha+2$ as $\alpha$ increases.
Figure 15 presents the tail latency and throughput of serving the OPT-13b model on the Alpaca dataset using the different layer selection strategies described above. Our findings show that using $m=\alpha+2$—where transferred layers share a two-layer memory region to enable double buffering—is necessary for optimal performance. Using $m=\alpha+1$ consistently (A) can result in a reduction in $12.7\%$ throughput compared to using $m=\alpha+2$ consistently (B) or switching to $m=\alpha+2$ as $\alpha$ increases (C). The reduction in the amount of data transfer when using $m=\alpha+1$ does not outweigh the overhead caused by disruptions to the GPU execution pipeline.
7.6. When and How Many Layers to Remap?
We conduct ablation studies to explore additional design parameter choices in MIRAGE. While parameter remapping needs to be initiated only when KV cache capacity is exhausted, determining whether to “halt” remapping during off-peak periods is also important (§ 7.6.1). Moreover, more aggressive remapping can help avoid recomputation and reduce tail latency, but it may increase average latency due to increased data transfer overhead (§ 7.6.2).
Figure 16. TBT CDF comparison between without and with Dynamic Reversion for OPT-13b+Alpaca. Dynamic Reversion effectively reduces TBT latency for non-peak periods.
7.6.1. When to halt remap?
MIRAGE monitors KV cache utilization at runtime and detects when free KV cache is available, typically during non-peak periods. For example, real-world traces show that request volume during non-peak times drops to approximately 20% of that seen during peak periods (Stojkovic et al., 2024b). To optimize performance in such periods, MIRAGE supports “halting” parameter remapping by reconfiguring the previously remapped memory space back for parameter usage. We name this mechanism Dynamic Reversion and evaluate it in this section.
Figure 16 shows the cumulative distribution function (CDF) comparison of TBT latency when serving the OPT-13b model on the Alpaca dataset, with and without the use of Dynamic Reversion. During peak load (e.g., arrival rate of 10.0 RPS), the Dynamic Reversion mechanism offers limited benefit, as parameter remapping is consistently needed to avoid recomputation. However, during non-peak periods (e.g., 1.0 RPS), Dynamic Reversion significantly improves performance, reducing P50 latency by 49% and consistently yielding lower TBT.
Figure 17. Capped vs non-capped remapping percentage. To prevent parameter loading from becoming the bottleneck, the remapping percentage must be appropriately capped.
7.6.2. How many layers to remap?
A key design parameter in MIRAGE is the remapping percentage—the proportion of model parameter memory dynamically repurposed for KV cache. Increasing the remapping percentage enables the system to support higher request volumes during peak periods, which are difficult to predict offline. However, aggressive remapping introduces higher CPU–GPU data transfer overhead, which can interfere with the performance of concurrent requests.
We evaluate the impact of remapping percentage using OPT-13b with Alpaca dataset. As shown in Figure 17, aggressive remapping reduces tail latency under light workloads (e.g., 1.0 RPS), but increases average latency due to higher CPU–GPU data transfer overhead. In contrast, capping the remapping percentage offers a more balanced performance: it sustains low average latency across a range of workloads while still reducing tail latency under variable request rates. For instance, at 10 requests per second, the capped strategy reduces P99 latency by 58% and P50 latency by 44%.
8. Discussion and Related Work
MIRAGE is highly compatible with existing LLM optimizations. Moreover, evolving hardware trends further emphasize the need for a system like MIRAGE to fully exploit the growing CPU-GPU bandwidth. We also highlight the key differences between MIRAGE and other related memory offloading or swapping approaches.
Compatibility with Other LLM Optimization Solutions: MIRAGE is designed to be highly compatible with state-of-the-art LLM inference optimizations. When the Dynamic Remapping Engine reaches its maximum remapping percentage, MIRAGE can further leverage CPU computation offloading-based methods, such as NEO (Jiang et al., 2024) and PowerInfer (Song et al., 2023), to ensure the quality of service for the online inference serving system. MIRAGE is also fully compatible with quantization methods (Lin et al., 2024; Zhao et al., 2024; Xia et al., 2024) and KV cache compression frameworks (Liu et al., 2023, 2024b, 2024a). Unlike KV cache swapping, MIRAGE requires no changes to CUDA kernels, making it compatible with advanced LLM kernels, such as Pod-Attention (Kamath et al., 2024) and Lserve (Yang et al., 2025). As a scheduler-independent memory management engine, MIRAGE supports both temporal and spatial sharing, making it compatible with a wide range of multi-tenant scheduling policies—including temporal (Bai et al., 2020; Gujarati et al., 2020; Xiao et al., 2020; Fu et al., 2024; Patke et al., 2024), spatial (Romero et al., 2021; Strati et al., 2024; Dhakal et al., 2020; Choi et al., 2022; Xu et al., 2022; Han et al., 2024a; Wang et al., 2024a), and hybrid approaches (Duan et al., 2024; Yu et al., 2025). Since MIRAGE activates parameter remapping only when GPU memory falls short for KV cache, it serves as a fallback mechanism for schedulers that mispredict memory demand under varied request arrival rates.
Evolving LLM Trends: The growing popularity of multi-agent workflows (Talebirad and Nadiri, 2023; Chan et al., 2023; Li et al., 2024b; Han et al., 2024b; Kim et al., 2025) makes MIRAGE particularly well-suited for scenarios involving temporal GPU resource sharing. Furthermore, as the trend toward deploying multiple smaller-scale LLMs on a single host becomes more prevalent (Wang et al., 2024b), MIRAGE offers additional benefits by enabling efficient memory reuse across co-located models. For emerging models with smaller KV caches enabled by multi-query attention (Dubey et al., 2024), MIRAGE remains compatible by repurposing the parameter memory of inactive models to expand the available memory for active ones—offering greater flexibility than KV cache swapping based approaches.
Evolving Hardware Trends: We believe the NVIDIA Grace Hopper Superchip is not just a special case but a clear demonstration of the growing trend toward increased CPU-GPU bandwidth in future hardware. NVLink, the underlying interconnect, is already mature and widely adopted. Its successor, GB200 (Nvidia, 2025b), preserves the 900 GB/s bidirectional bandwidth while scaling to 2 GPUs per node and up to 72 GPUs in an NVLink-connected rack-scale system. This architecture ensures that MIRAGE remains compatible with multi-GPU and distributed LLM inference optimizations (Sun et al., 2024; Fu et al., 2024; Yu et al., 2025). The growth in CPU performance, along with ongoing improvements in CPU DRAM bandwidth, further reinforces the trend toward high CPU–GPU bandwidth. Unlike traditional DDR memory, the LPDDRX memory used in GH200 offers significantly higher bandwidth with lower power consumption, enabling such high CPU-GPU transfer rates (Lee et al., 2018; Park et al., 2024; Li et al., 2018). As memory subsystems become increasingly critical in datacenter architectures (Maruf et al., 2023; Mahar et al., 2023), LPDDRX adoption is expected to grow, further supporting this trend.
Related Works: Although prior work has explored using CPU memory as an extension of GPU memory, the focus has predominantly been on training scenarios, as seen in systems such as DeepSpeed (Rajbhandari et al., 2020; Yao et al., 2024; Ren et al., 2021; Rajbhandari et al., 2021) and others (Huang et al., 2020; Wang et al., 2022; Yuan et al., 2024; Lim et al., 2021). Training workloads are generally compute-bound, making them more tolerant of memory offloading overhead. In contrast, inference serving is latency-sensitive rather than throughput-oriented, making such memory offloading significantly more challenging. Consequently, only a few recent efforts have explored memory swapping in the context of LLM inference serving. Frameworks like FlexGen (Sheng et al., 2023) perform on-demand data migration from CPU to GPU, which can introduce delays as computations are forced to wait for data transfers to finish (Xu et al., 2024). Pie (Xu et al., 2024) was the first to explore the benefits of increased CPU-GPU bandwidth by swapping KV cache. However, its performance is affected by the bidirectional data transfer. MIRAGE stands apart from all prior work also as the first to support multi-tenant LLM inference serving.
9. Conclusion
This paper introduces MIRAGE, a Dynamic Remapping Engine that automatically and elastically repurposes model parameter memory to expand KV cache capacity for LLM inference serving, enabling low tail latency and high throughput during runtime. MIRAGE leverages the substantially increased CPU–GPU bandwidth in modern systems to overlap parameter transfers with GPU computation, minimizing the “remapping” overhead. MIRAGE can be deployed in any serving system and remains compatible with a wide range of popular multi-tenant GPU sharing and scheduling strategies. MIRAGE achieves 6.6%–86.7% improvement in throughput and 20.7%–99.3% reduction in tail latency compared to vLLM. Additionally, compared to KV cache swapping, parameter remapping incurs lower data transfer and synchronization overhead while offering greater flexibility in multi-tenant environments.